# theorem]Construction

### v1

@@ 6.1. Trees
Let us first see that if a $T$-process goes on for t2t-2 rounds then it stabilises in the next round. Observe that the centre vertex $c_{i}$ of a copy of $K_{1,t-1}$ created in the $i^{\text{th}}$ round for some $i\leq t-2$, becomes universal at round $i$. This means that after round t2t-2 every vertex has at least t2t-2 neighbours. So all non-universal vertices are now centres of $K_{1,t-1}$ with one edge missing, and in the next round all vertices become universal and the $T$-process stabilises by percolating.
@@ 7. Graph parameters and the running time
In Theorem 6.3 and Theorem 6.6 we saw that the maximum running time of the $H$-process is constant for trees and logarithmic for cycles, respectively. These are in stark contrast with cliques of order $4$, where the running time of the process could be at least linear, and the wide array of graphs encountered in Section 5 where the running time could be (almost) quadratic. Which properties of the infection rule $H$ cause the $H$-process to always finish quickly? This is what we study in this section, and try to identify necessary and sufficient conditions for sub-linear or linear maximum running time, in terms of relevant graph parameters, involving the degrees, the connectivity and the tree-width of the infection rule. In this section we focus on connected infection rules (see Section 9 for the disconnected case).
@@ 7.1. The role of the degrees
In our analysis of trees, the existence of leaves plays a crucial role in showing that the running time is constant. In the analysis of the percolation threshold $p_{c}(H)$ [12] in $G(n,p)$ the presence of a single pendant vertex in the infection rule $H$ also plays a central role in pushing the threshold towards a sparser range (see Section 4.2). One cannot help but wonder whether the minimum degree being $1$ has anything to do with small maximum running time in general. It turns out that, with the exception of cycles, it is indeed necessary for a sub-linear running time.
@@
This implies that in our quest towards a characterisation of connected graphs $H$ with sub-linear running time we need to identify additional conditions besides the minimum degree being $1$. Can limiting the maximum degree help? The graph $H$ in Theorem 7.4 has a vertex of full degree (of degree $15$, see Figure 5). The construction in our next result shows that even limiting the maximum degree to $3$ is not helpful to guarantee sub-linear maximum running time.
@@
It is possible to slightly modify the proofs of the above Theorems 7.4 and 7.5 to show that an additional condition in terms of the average degree of $H$ is of no help in guaranteeing faster running time. This modification of Theorem 7.5 for example creates infection rules with minimum degree $1$, average degree arbitrarily close to $2$, and yet, quadratic maximum running time. On the other hand, connected infection rules with average degree at most $2$ are unicyclic. These are trees with exactly one extra edge. It is possible, by a small modification of the proof of Theorem 6.3, to show that these, with the exception of cycles, have constant maximum running time.
@@ 7.2. Proof technique: Simulation
A good plan, a nice plan, easier to dream up though than to actually carry it out. The first idea would be just to attach a pendant edge to $K_{6}$, but this was shot down already in Example 7.3 to have maximum running time $3$. The main problem with adding a pendant vertex to a “slow infection rule” becomes highly visible: the pendant edge is added immediately to any vertex in any copy of $K_{6}$, creating many universal vertices and a speedy finish.
@@
Figure 5. A graph $H$ with minimum degree $1$ and quadratic running time.
The graph $H$ consists of three pairwise vertex-disjoint parts: A clique $B$ of size six, a clique $A$ of size nine with a special vertex $u$ which is universal, and a pendant vertex $v^{*}$ attached to it. As the starting graph $\tilde{G}$ we take the starting graph $G$ of the slow $K_{6}$-process defined by Balogh et al. [14], together with a disjoint copy $K$ of $K_{9}$, and make one of its vertices $w$ fully connected to $V:=V(G)$. Then any copy of $K_{6}-e$ in $V$ during the $H$-process, together with $K$ and an appropriate pendant edge from $w$ back to $V$ forms a copy of HeH-e. So all edges that get infected in the $K_{6}$-process on $G$ will also get infected in the $H$-process on $\tilde{G}$. We still need to make sure though, that the above described are all the copies of HeH-e during the $H$-process on $\tilde{G}$ and hence the process does not speed up compared to the $K_{6}$-process on $G$, due to an unexpected copy of $H$ minus one of its edges somewhere. The analysis of the $K_{6}$-process $(G_{i})_{i\geq 0}$ given in [14] implies that no copy of $K_{7}$ appears for $\Omega(n^{2})$ rounds. Using this and an induction over the rounds, one can show that in any copy of HeH-e the set $A$ has to be mapped to $K$ and $u$ to $w$, as required.
The basic idea of this simulation construction method is flexible and can be adapted to different settings. The two main ingredients are some benchmark infection rule $B$ (this was $K_{6}$ above) whose process and running time we intend to simulate and an appropriate anchor graph $A$ (this was $K_{9}$ above) to which we can connect the benchmark graph $B$ as well as pendent edges to construct $H$. The starting graph will then be the starting graph for $B$ and a copy of $A$ attached to it. The aim is to show that $A$ can only appear as itself in the copies of $H$ of the $H$-process so unplanned copies of HeH-e do not appear. To achieve this we need to construct an anchor graph which is “far enough” from anything that will ever appear in the simulated $B$-process on $G$.
For example to prove Theorem 7.5, these contemplations motivate us towards the construction of the graph depicted in Figure 6. The benchmark infection rule $B$ is $C_{6}$ with a chord which has at least linear running time by Theorem 7.1. The anchor graph $A$ is a more complex ad hoc construction, together with carefully chosen attachment vertices $u$ and $u^{\prime}$ (which are now different from each other to control the maximum degree). The proof that this construction works hinges on an extension [46] of Theorem 7.1 that guarantees that there is some $B$-process $(G_{i})_{i\geq 0}$ with $\tau_{B}(G_{0})$ linear, such that all of the $G_{i}$ are bipartite. The crucial property of the anchor graph is that it contains many triangles, and so is far from bipartite.
@@ 7.5. A conjecture on tree-width
Given Theorem 6.3, another natural parameter to consider is the tree-width ${\mathrm{tw}}(H)$ of the infection rule $H$. It is well known that ${\mathrm{tw}}(H)\geq\delta(H)\geq\kappa(H)$. Example 7.3 shows that unlike small minimum degree and small connectivity, small tree-width is not necessary for small maximum running time, not even constant. Indeed, for the infection rule $K_{k}^{+}$ with maximum running time $3$ we have ${\mathrm{tw}}(K_{k}^{+})={\mathrm{tw}}(K_{k})=k-1$. However, it might still be the case that small tree-width is sufficient to imply an upper bound on maximum running time. Indeed, trees are the unique graphs $H$ with ${\mathrm{tw}}(H)=1$ and thus Theorem 6.3 is simply saying that any graph with tree-width 1 has constant running time. Graphs of tree-width 2 can already have linear running time, by considering the complete bipartite graphs $K_{2,s}$ and noting that ${\mathrm{tw}}(K_{k,\ell})=\min\{k,\ell\}$.
@@ 8. Bipartite infection rules
Let $(G_{i})_{i\geq 0}$ be an $H$-process with $\tau=\tau_{H}(G_{0})=M_{H}(n)\geq 2$. For $1\leq i\leq\tau$, let $e_{i}\in G_{i}\setminus G_{i-1}$ be an edge added at time $i$. Now consider the graph $G^{\prime}$ with
@@
Indeed, for many graphs, like trees and unicyclic graphs, we already know from Theorems 6.3 and 6.6 that this upper bound is not tight, by far in fact. For complete bipartite graphs with a part of size two we can also do much better, as their maximum running time is linear by Proposition 7.13, while the Turán number is of order $n^{3/2}$. Before completely disrespecting Proposition 8.1 though, we should reflect on what happens for non-bipartite graphs. The generous upper bound is tight [14] up to a constant factor for cliques of order at least $6$, and in fact also for “most” graphs by Theorem 5.3. So maybe we are also in for a surprise for large enough bipartite $H$?
@@ 8.1. Small graphs
The bound on the minimum degrees is best possible. Indeed, for an arbitrary $k$ divisible by $4$ consider two copies of $K_{k/4,k/4}$ connected with a single edge. This graph is $(2,1)$-separable and self-percolating and hence has at most linear running time by Theorem 7.10. The minimum degree on both sides of the bipartition is $k/4$, exactly half of the sides.
@@
One could use the slow $K_{3,3}$-process in the simulation construction method (Section 7.2) to create graphs with minimum degree $1$, maximum degree $4$ and maximum running time $n^{3/2}$. In the proof of Theorem 7.5 one simply replaces the benchmark graph ($C_{6}$ plus a chord) with $K_{3,3}$. The obtained infection rule has a single vertex of degree $4$. In order to get down to maximum degree $3$, one needs to delete an edge from $K_{3,3}$; unfortunately this reduces the maximum running time to linear. One can also aim to create a larger almost $3$-regular bipartite infection rule randomly and hope that it has a longer running time. We are not convinced though that such a graph would have super-linear running time either.
@@ 9. Disconnected infection rules
9.1. Forests and $2$-regular graphs
@@
As for unions of cycles, i.e. $2$-regular graphs, recall that when the infection rule is $C_{k}$ any two vertices of distance k1k-1 become adjacent in each round of the process, which led to the maximum running time on $n$-vertex graphs essentially being $\log_{k-1}(n)$. In the case that $H$ is the disjoint union of cycles the maximum running time is controlled by the cycle that provides the fastest reduction of the diameter, which is just the largest cycle in $H$.
@@ 9.3. Interacting infection rules
This example further tells us that Corollary 7.2 breaks down for disconnected infection rules. We remark though that Theorem 7.1 can be extended to disconnected $H$ by the stronger requirement that every component of $H$ has minimum degree at least $2$ and maximum degree at least $3$.

### v2

@@ 6.1. Trees
Let us first see that if a $T$-process goes on for t2t-2 rounds then it stabilises in the next round. Observe that the centre vertex $c_{i}$ of a copy of $K_{1,t-1}$ created in the $i^{\text{th}}$ round for some $i\leq t-2$, becomes universal at round ii. This means that after round t2t-2 every vertex has at least t2t-2 neighbours. So all non-universal vertices are now centres of $K_{1,t-1}$ with one edge missing, and in the next round all vertices become universal and the $T$-process stabilises by percolating.
@@ 7. Graph parameters and the running time
In Theorem 6.3 and Theorem 6.6 we saw that the maximum running time of the $H$-process is constant for trees and logarithmic for cycles, respectively. These are in stark contrast with cliques of order 44, where the running time of the process could be at least linear, and the wide array of graphs encountered in Section 5 where the running time could be (almost) quadratic. Which properties of the infection rule $H$ cause the $H$-process to always finish quickly? This is what we study in this section, and try to identify necessary and sufficient conditions for sub-linear or linear maximum running time, in terms of relevant graph parameters, involving the degrees, the connectivity and the tree-width of the infection rule. In this section we focus on connected infection rules (see Section 9 for the disconnected case).
@@ 7.1. The role of the degrees
In our analysis of trees, the existence of leaves plays a crucial role in showing that the running time is constant. In the analysis of the percolation threshold $p_{c}(H)$ [12] in $G(n,p)$ the presence of a single pendant vertex in the infection rule $H$ also plays a central role in pushing the threshold towards a sparser range (see Section 4.2). One cannot help but wonder whether the minimum degree being 11 has anything to do with small maximum running time in general. It turns out that, with the exception of cycles, it is indeed necessary for a sub-linear running time.
@@
This implies that in our quest towards a characterisation of connected graphs $H$ with sub-linear running time we need to identify additional conditions besides the minimum degree being 11. Can limiting the maximum degree help? The graph $H$ in Theorem 7.4 has a vertex of full degree (of degree $15$, see Figure 5). The construction in our next result shows that even limiting the maximum degree to 33 is not helpful to guarantee sub-linear maximum running time.
@@
It is possible to slightly modify the proofs of the above Theorems 7.4 and 7.5 to show that an additional condition in terms of the average degree of $H$ is of no help in guaranteeing faster running time. This modification of Theorem 7.5 for example creates infection rules with minimum degree 11, average degree arbitrarily close to 22, and yet, quadratic maximum running time. On the other hand, connected infection rules with average degree at most 22 are unicyclic. These are trees with exactly one extra edge. It is possible, by a small modification of the proof of Theorem 6.3, to show that these, with the exception of cycles, have constant maximum running time.
@@ 7.2. Proof technique: Simulation
A good plan, a nice plan, easier to dream up though than to actually carry it out. The first idea would be just to attach a pendant edge to $K_{6}$, but this was shot down already in Example 7.3 to have maximum running time 33. The main problem with adding a pendant vertex to a “slow infection rule” becomes highly visible: the pendant edge is added immediately to any vertex in any copy of $K_{6}$, creating many universal vertices and a speedy finish.
@@
Figure 5. A graph $H$ with minimum degree 11 and quadratic running time.
The graph $H$ consists of three pairwise vertex-disjoint parts: A clique $B$ of size six, a clique AA of size nine with a special vertex $u$ which is universal, and a pendant vertex $v^{*}$ attached to it. As the starting graph $\tilde{G}$ we take the starting graph $G$ of the slow $K_{6}$-process defined by Balogh et al. [14], together with a disjoint copy $K$ of $K_{9}$, and make one of its vertices $w$ fully connected to $V:=V(G)$. Then any copy of $K_{6}-e$ in $V$ during the $H$-process, together with $K$ and an appropriate pendant edge from $w$ back to $V$ forms a copy of HeH-e. So all edges that get infected in the $K_{6}$-process on $G$ will also get infected in the $H$-process on $\tilde{G}$. We still need to make sure though, that the above described are all the copies of HeH-e during the $H$-process on $\tilde{G}$ and hence the process does not speed up compared to the $K_{6}$-process on $G$, due to an unexpected copy of $H$ minus one of its edges somewhere. The analysis of the $K_{6}$-process $(G_{i})_{i\geq 0}$ given in [14] implies that no copy of $K_{7}$ appears for $\Omega(n^{2})$ rounds. Using this and an induction over the rounds, one can show that in any copy of HeH-e the set AA has to be mapped to $K$ and $u$ to $w$, as required.
The basic idea of this simulation construction method is flexible and can be adapted to different settings. The two main ingredients are some benchmark infection rule $B$ (this was $K_{6}$ above) whose process and running time we intend to simulate and an appropriate anchor graph AA (this was $K_{9}$ above) to which we can connect the benchmark graph $B$ as well as pendent edges to construct $H$. The starting graph will then be the starting graph for $B$ and a copy of AA attached to it. The aim is to show that AA can only appear as itself in the copies of $H$ of the $H$-process so unplanned copies of HeH-e do not appear. To achieve this we need to construct an anchor graph which is “far enough” from anything that will ever appear in the simulated $B$-process on $G$.
For example to prove Theorem 7.5, these contemplations motivate us towards the construction of the graph depicted in Figure 6. The benchmark infection rule $B$ is $C_{6}$ with a chord which has at least linear running time by Theorem 7.1. The anchor graph AA is a more complex ad hoc construction, together with carefully chosen attachment vertices $u$ and $u^{\prime}$ (which are now different from each other to control the maximum degree). The proof that this construction works hinges on an extension [46] of Theorem 7.1 that guarantees that there is some $B$-process $(G_{i})_{i\geq 0}$ with $\tau_{B}(G_{0})$ linear, such that all of the $G_{i}$ are bipartite. The crucial property of the anchor graph is that it contains many triangles, and so is far from bipartite.
@@ 7.5. A conjecture on tree-width
Given Theorem 6.3, another natural parameter to consider is the tree-width ${\mathrm{tw}}(H)$ of the infection rule $H$. It is well known that ${\mathrm{tw}}(H)\geq\delta(H)\geq\kappa(H)$. Example 7.3 shows that unlike small minimum degree and small connectivity, small tree-width is not necessary for small maximum running time, not even constant. Indeed, for the infection rule $K_{k}^{+}$ with maximum running time 33 we have ${\mathrm{tw}}(K_{k}^{+})={\mathrm{tw}}(K_{k})=k-1$. However, it might still be the case that small tree-width is sufficient to imply an upper bound on maximum running time. Indeed, trees are the unique graphs $H$ with ${\mathrm{tw}}(H)=1$ and thus Theorem 6.3 is simply saying that any graph with tree-width 1 has constant running time. Graphs of tree-width 2 can already have linear running time, by considering the complete bipartite graphs $K_{2,s}$ and noting that ${\mathrm{tw}}(K_{k,\ell})=\min\{k,\ell\}$.
@@ 8. Bipartite infection rules
Let $(G_{i})_{i\geq 0}$ be an $H$-process with $\tau=\tau_{H}(G_{0})=M_{H}(n)\geq 2$. For $1\leq i\leq\tau$, let $e_{i}\in G_{i}\setminus G_{i-1}$ be an edge added at time ii. Now consider the graph $G^{\prime}$ with
@@
Indeed, for many graphs, like trees and unicyclic graphs, we already know from Theorems 6.3 and 6.6 that this upper bound is not tight, by far in fact. For complete bipartite graphs with a part of size two we can also do much better, as their maximum running time is linear by Proposition 7.13, while the Turán number is of order $n^{3/2}$. Before completely disrespecting Proposition 8.1 though, we should reflect on what happens for non-bipartite graphs. The generous upper bound is tight [14] up to a constant factor for cliques of order at least 66, and in fact also for “most” graphs by Theorem 5.3. So maybe we are also in for a surprise for large enough bipartite $H$?
@@ 8.1. Small graphs
The bound on the minimum degrees is best possible. Indeed, for an arbitrary $k$ divisible by 44 consider two copies of $K_{k/4,k/4}$ connected with a single edge. This graph is $(2,1)$-separable and self-percolating and hence has at most linear running time by Theorem 7.10. The minimum degree on both sides of the bipartition is $k/4$, exactly half of the sides.
@@
One could use the slow $K_{3,3}$-process in the simulation construction method (Section 7.2) to create graphs with minimum degree 11, maximum degree 44 and maximum running time $n^{3/2}$. In the proof of Theorem 7.5 one simply replaces the benchmark graph ($C_{6}$ plus a chord) with $K_{3,3}$. The obtained infection rule has a single vertex of degree 44. In order to get down to maximum degree 33, one needs to delete an edge from $K_{3,3}$; unfortunately this reduces the maximum running time to linear. One can also aim to create a larger almost 33-regular bipartite infection rule randomly and hope that it has a longer running time. We are not convinced though that such a graph would have super-linear running time either.
@@ 9. Disconnected infection rules
9.1. Forests and 22-regular graphs
@@
As for unions of cycles, i.e. 22-regular graphs, recall that when the infection rule is $C_{k}$ any two vertices of distance k1k-1 become adjacent in each round of the process, which led to the maximum running time on $n$-vertex graphs essentially being $\log_{k-1}(n)$. In the case that $H$ is the disjoint union of cycles the maximum running time is controlled by the cycle that provides the fastest reduction of the diameter, which is just the largest cycle in $H$.
@@ 9.3. Interacting infection rules
This example further tells us that Corollary 7.2 breaks down for disconnected infection rules. We remark though that Theorem 7.1 can be extended to disconnected $H$ by the stronger requirement that every component of $H$ has minimum degree at least 22 and maximum degree at least 33.